Sol Summary – February 8th

Crew 204 Sol Summary Report 08-FEB -2019

Sol: 11

Summary Title: The End Game

Author’s name: Kunal Naik (Space_Bot)

Mission Status: Nominal

Sol Activity Summary: Time to go home. Is it happiness, eagerness or desperation to see actual beings around us, The Psychological game is going to end and what we learn in 2 weeks? Was it something changing?

Yes, of course not a life changing one but surely a lesson to live on limited water, food supply and strict protocols.

Experiencing some percentage of daily life of Astronauts. The first thing which the crew realized after waking up is the final day of the sim.

The crew had the morning breakfast and were in a rush for their morning EVA, Carbon Flight and MarsBag went for EVA while Cosmos and SpaceBot were in their daily Hab activities. Today’s in charge for cooking was SpaceBot, Post EVA, the crew had lunch together with challenges of Sim as a discussion topic. Today there were 2 EVA’s, one in the morning while the second after lunch. This time, it was SpaceBot and Cosmos turn for the EVA. While other crew members rested for a while at the Hab. Green Positive vibes all over the Green Hab. It’s sad that the crew won’t be able to see the growth of their planted seeds. This is the Ending and probably the last Sol Summary.

Look Ahead Plan: The crew has planned to clean the entire facility from top to bottom, from right to left. The crew will be eagerly waiting to welcome, meet and handing over the HAB to the Crew 205 for future simulation.

Anomalies in work: None Encountered

Weather: Sunny

Crew Physical Status: Nominal

EVA’s location: 1) 516100 E; 4254550 N (Widows Peak) 2) 518250 E;
4255500 N (Quarry Road)

Reports to be file: 1) Operations Report 2) Green Hab Report 3)
Journalist Report 4) EVA Report 5) Photos 6) Mission Summary 7) End
Rotation Food Supply excel sheet

Support Requested: None

Greenhab Report – February 8th

Crew 204 Green Hab Report – 8-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 16.3°C

High: 27.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 171 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 19.4° C / 27 %
2) 12.00 PM 27.0° C / 20 %
3) 15:00 PM 18.0° C / 22 %
4) 18:00 PM 16.3° C / 32 %

Changes to crops: None

Narrative: Out of all the sprouts, fenugreek sprouts have grown faster. Cover shaped leaves are about 5 cm high till today and growing, which make a pretty groundcover between taller, widely spaced crops. and the fact that fenugreek roots are capable of fixing nitrogen in the soil to naturally fertilize the crops around them. All these make the crop a perfect all-rounder.

Harvest: None

Support/supplies needed: None

Operations Report – February 8th

Crew 204 Operations Report 08- FEB – 2019

SOL: 11

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1830 (8th Feb), Stopped by this morning 0900 (8th Feb)

Solar— SOC % – Turned off Gen (8th Feb) 77%, Turned on Gen (8th Feb) 72%

Diesel Reading – 30%

Propane Reading – ~ 55%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 58%, ~ 290 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 168 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01409242 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: 1st EVA

Hours: 48.5 hours

Beginning charge: 100%

Ending charge: 42%

Currently charging: Yes

Curiosity rover used: 1st EVA

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 12%

Currently charging: Yes

Spirit rover used: 2nd EVA

Hours: 72.8

Beginning charge: 100%

Ending charge: 25%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? The car is in Workshop, not at the campus

Crew Car used and why, where? Off Campus

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Journalist Report – February 8th

Crew 204 Crew Journalist Report 08-FEB-2019

Author: Sonal Baberwal

Title: When we left Earth!

Good Morning Earthlings! The final morning on this soil. Yes, we are returning back to the Earth. Probably the re-entry module has been aligned on the launchpad. But we will miss this place a lot. A completely new world! This isolation acted as a detox from the regular world. It was necessary to have self-analysis. A lot of things have been learnt from this mission.

Another morning starts where the crew had been involved in cleaning the Hab. we had the final two general purpose EVAs. the morning EVA was performed by the Commander and myself followed by the GreenHab Officer and the Crew Engineer.

The day passed by really quickly with no awareness of time! I would like to share a short poem I wrote for MDRS experience, pardon me if not up to date!

“The plants, the lab,
The rovers, the Hab.
The love, the care,
The friendship we share.
The food to cook,
The windows we look.
The signing in and signing off,
The mission support and Cap Com.
The EVAs we risked,
Yes, everything will be missed!”
~ Sonal Baberwal.

Congratulations to the Crew and the mission support for this successful mission. This is Sonal Baberwal, HSO and Crew Journalist reporting live for the last time from the upper deck of the Habitat and signing off on the final day! I wish very good luck and best wishes to Crew 205. Stay connected to hear new stories from Crew 205!

Operations Report- February 7th

Crew 204 Operations Report 07- FEB – 2019

SOL: 10

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1922 (7th Feb),
Stopped by this morning 0630 (7th Feb)

Solar— SOC % – Turned off Gen (7th Feb) Not checked, Turned on Gen
(7th Feb) 77%

Diesel Reading – 32%

Propane Reading – ~ 58%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 60%, ~ 300 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 183 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01408827 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Spirit rover used: Not Used
Hours: 72.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: Not Used
Hours: 47.6 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Not Used
Hours: 72.7 hours

Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? The car is in Workshop

Crew Car used and why, where? Off Campus

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Greenhan Report- February 7th

Crew 204 Green Hab Report – 7-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.2°C

High: 26.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 183 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 22.2° C / 39 %
2) 12.00 PM 24.9° C / 21 %
3) 15:00 PM 19.9° C / 21 %
4) 18:00 PM 18.2° C / 26 %

Changes to crops: None

Narrative: After almost 12 gallons of water, hours of supplementary light and lots of care later, plants are noted to be in great shape and health.

Harvest: None

Support/supplies needed: None

EVA Report- February 7th

Crew 204 EVA Report 07-FEB-2019

EVA # 8

Author: Avishek Ghosh.

Purpose of EVA: A short duration EVA to demonstrate how to operate rover remotely.

Start time: 08:00 hrs.

End time: 08:45 hrs.

Narrative: EVA started on time with a bright morning under a moderate wind. EVA crew were able to communicate remotely with the rover group from Sipna Engineering college in India. The group was able to control the child rover remotely over wi-fi and performed a couple of iterations. Probably this is the first time a rover is controlled and operated remotely (around 13,000 km far) on MDRS terrains.

Destination: Surrounding of Hab; mostly close to both sides of the Greenhab.

Coordinates: 518100 E ; 4250750 N

Participants: Kunal and Sonal.

Roads and routes per MDRS map: Close to hab (Route 1103).

Mode of travel: Walking.

Journalist Report- February 7th

Crew 204 Journalist Report 07-FEB-2019

Author: Sonal Baberwal (HSO & Crew Journalist)

Good Morning from Mars!

The morning seems early today. This was the day we all have been eagerly waiting for! A lot of hope and enthusiasm had been preserved for this day. Today the crew woke up pretty early than the regular schedule as the day had something special. We all started preparing for the EVA.

This has been really exciting because our rover is ready and scheduled to be operated from Earth (India). We got ready and suited up in the Spacesuit with heartbeats that could be heard loud. Will this attempt be a successful one? Is that a dream coming true that a rover will be operated from a remote location from Earth on MDRS terrains? Yes, it is possible! Salute to our technology and emerging IoT trend!

The rover had been operated for 51 seconds from Earth by Mr. Harish Ajankar from Sipna College of Engineering and Technology, India in front of the Professors of the University. The biggest challenge was to find a flexible time for the EVA that supports different time zones. Thanks to mission support for their cooperation and support to make this happen into reality.

It was a pretty good experience to observe the rover crawling on the MDRS terrains which have been operated from India. I believe this legacy will turn into an inspirational and motivational example for the young generation and someday participate in this simulation.

The day has been followed by daily activities. The evenings have been really good as we have been enjoying games and movies together living the final days of simulation to the fullest!

The day had been really interesting in terms of the activities we have been performing. Looking forward to a new sunrise. This is Sonal Baberwal, HSO and Crew Journalist reporting live from the upper deck of the Habitat! Stay connected to know more about our last days of the simulation.

Sol Summary- February 7th

Crew 204 Sol Summary Report 07- FEB – 2019

Sol: 10

Summary Title: The 51 second excitation

Author’s name: Kunal Naik (Space_Bot)

Mission Status: Nominal

Sol Activity Summary: What gives an excitation? A rover working from ~ 13000 km with live Audio/Video stream via rover will surely create one. Today was the judgment day for the child rover project. The crew woke up earlier than their assigned schedule, potato creamy soup is a good start of the day. The crew was on the run as the EVA was requested post breakfast.

SpaceBot and MarsBag carried out the walking EVA near the MDRS HAB premises. The main purpose of the EVA was to operate child rover from a remote location, India (13000 kms away). It was operated by students of Sipna College of Engineering. The demonstration was successful in exploring the Martian soil. The crew was happy as their project was a success. Still needs upgradation for sure.

Post EVA, the team discussed about the rover and Indian Mythology. Pre-Lunch, the crew members were on their assigned respective task. Lunch was prepared by CarbonFlight and SpaceBot, quite a lot food which will be used for dinner as well.

Post lunch, all crew members rested for a while. Carbon flight printed few bricks
for Martian habitat from his 3D Clay Printer a day ago and now they are dried and ready to be baked. There is a positive vibe from the Green Hab, every-day you eager to go to the GreenHab just to see the growth of the plants. The crew is still planning for some board games
post dinner.

Look Ahead Plan: The 3D clay printer has printed some bricks out of Martian soil composition which will be baked to test the inner dynamic structure and possibility to use the brick for future habitat. The End is 2 days short, the crew has scheduled the recleaning of the hab.

The sprouts and mix seeds have already started showing good results. Hoping to observe more growth.

Anomalies in work: None Encountered

Weather: Sunny

Crew Physical Status: Nominal

EVA’s location: Hab Premise

Reports to be file: 1) Operations Report 2) Green Hab Report 3)
Journalist Report 4) EVA Report 5) Photos

Support Requested: None